### v2

# Linear analysis of I-C-Love universal relations for neutron stars

Abstract
Neutron stars (NSs) are excellent laboratories for testing gravity theories as they are strongly self-gravitating bodies and have rich observational phenomena. However, strong-field gravity effects in NS could be degenerate with their equation of state (EOS) which is largely unknown. Fortunately, there exist the so-called universal relations among the NS macroscopic quantities that are found to be insensitive to the underlying EOS. Studying the origin of these relations can lead to a better understanding of NSs and the gravitational interaction. We develop a new perspective of view to analyze the I-C and I-Love universal relations for NSs. At the linear order, we separate the deviation of the universal relations into two factors that are multiplied together. One is the EOS difference while the other factor only depends on the background star structure. The smallness of the second factor that does not depend on the EOS difference then indicates the origin of the universality. We discuss the validity of our linear approximation when considering the difference among realistic EOSs. Our study can be regarded as a new frame for quantitative representation of the universality and may provide new insights to the universal relations of NSs.
I Introduction
Neutron stars (NSs) are extremely compact objects in the Nature. Thanks to their strong self-gravity, large density in the inner region, and abundant observational phenomena, they are ideal laboratories for studying fundamental physics, including the strong-field gravity and nuclear physics Lattimer and Prakash (2001); Shao (2023); Freire and Wex (2024). For example, the mass-radius relationship of NSs strongly depends on the supranuclear matter equation of state (EOS), which is the relation between the matter density $\rho$ and pressure $p$. Thus, independently measuring the mass $M$ and radius $R$ of a NS can provide valuable constraints on the EOS at the high density regime that is unachievable with terrestrial nuclear experiments. The observations of heavy pulsars with masses $\gtrsim 2\,M_{\odot}$ have already ruled out those EOSs that are too “soft” for supporting a large mass NS in the general relativity (GR) Fonseca and others (2021); Demorest et al. (2010); Antoniadis and others (2013). Current X-ray observations with the NICER satellite can measure the mass and radius of a pulsar to $\sim 10\%$ accuracy and have also been used to constrain the EOS Riley and others (2019); Raaijmakers and others (2019).
However, the mass-radius relationship of NSs can be affected by the underlying gravity theory as well. A well-known example is the scalar-tensor theory, while passing all weak-field gravity tests in some parameter space Will (2018), predicts large deviation in the strong-field regime with a phenomenon called spontaneous scalarization Damour and Esposito-Farèse (1993, 1996); Doneva et al. (2024). Binary NSs were used as excellent testbeds, notably with the observational absence of the gravitational dipole radiation that is otherwise predicted in the scalar-tensor gravity theory Freire et al. (2012); Shao et al. (2017); Zhao et al. (2022). But still, considering a massive scalar field can evade these dipole-radiation constraints Ramazanoğlu and Pretorius (2016); Doneva and Yazadjiev (2016); Hu et al. (2020); Xu et al. (2020); Hu et al. (2021).
The degeneracy between the strong-field gravity effects and the nuclear matter EOS in NSs causes a challenge for studying both aspects Shao (2019). A test of strong-field gravity can only be done if we assume that we know the underlying EOS or systematically taken into account its large uncertainties, and vise versa. One way to break this degeneracy is provided by the finding of the so-called universal relations among some NS properties Yagi and Yunes (2013b, a); Pappas and Apostolatos (2014); Chakrabarti et al. (2014); Gao et al. (2023); Luk and Lin (2018); Shao and Yagi (2022). These relations, for example, the well-known I-Love-Q relation for the dimensionless NS moment of inertia $\tilde{I}$, tidal Love number $\tilde{\Lambda}$ and spin induced quadrupole moment $\tilde{Q}$, are largely insensitive to the EOSs but sometimes do depend on the gravity theories or even not universal in theories beyond GR Yagi and Yunes (2017). Thus, independent measurement of these quantities will enable us to have EOS independent test of gravity theories. In this regard, the I-Love relation will be especially interesting as the tidal Love number of NS can be measured with gravitational wave observations Abbott and others (2018), while it is expected to measure the NS moment of inertia with the timing observation of the Double Pulsar J0737-3039A/B Hu et al. (2020); Kramer and others (2021).
Besides the interesting applications of these universal relations, studying themselves also provides us better understanding of NSs and gravity theories. The origin of the universality, which can be as good as $\lesssim 1\%$ level in the I-Love-Q relations, is still in discussion. In literature, it is proposed at the beginning that the universality may come from the similarity of the different EOSs in the low density regime or it is the remnants of the no-hair theorem Yagi and Yunes (2013a, b). Considering that some relations like I-Love-Q even hold among NSs and quark stars (QSs), which are composed with quark matter that has a completely different EOS behavior in the low density regime, it is then proposed that the universality may come from the emergent self-similarity of isodensity contours inside NSs Yagi et al. (2014) or the realistic EOS is “close” to the incompressible limit Sham et al. (2015); Katagiri et al. (2025). There are also analytical calculations of these relations based on series expansion of the star’s compactness and assuming an incompressible EOS Chan et al. (2015) or Tolman VII model that phenomenologically describes the density profile inside NSs Jiang and Yagi (2020); Lowrey et al. (2025).
Most of the studies on analyzing the NS universal relations take some specific parametrization of the EOS or assumptions of the NS configurations, and the relations are also often expanded around the Newtonian limit. These attempts may reveal the possible key ingredient that preserves the universality, but also may lead to questions that why these configurations hold for various EOSs. Though there are a limited number of studies using different approaches, Chan et al. (2016) showed that the I-Love relation in the incompressible limit is stationary at the leading order of EOS variation and star’s compactness.
In this work, we analyze the NSs’ I-C and I-Love universal relations from a new perspective of view. Rather than parametrizing or expanding the EOSs or NS structures, we consider the linear response of the universal relation under any EOS perturbations. With this approach, we are able to separate the deviation of the universal relation caused by EOS perturbations into two parts that are multiplied together: one is the difference in the EOSs and the other factor only depends on the background star solution. As one can have almost arbitrary EOS difference if the universal relation holds, the smallness of the second factor fully accounts for the universality. Our approach also does not assume an expansion in the star’s compactness thus is applicable for both Newtonian and GR configurations.
The remaining part of this paper is organized as follows. In Sec. II we display the basic equations for calculating the I-C universal relations, and we introduce our concepts and approaches in Sec. III. We apply our approach to a simple case that can be solved analytically in Sec. IV, and we apply the approach for realistic EOSs in Sec. V. Section VI shows the results for I-Love relation. We also discuss the validity of our linear approximation when considering the difference between realistic EOSs in Sec. VII. Finally, we summarize our work in Sec. VIII. Throughout this work, we use the geometric units of $G=c=1$.
II I-C universal relation
Here we briefly review the I-C universal relation and the basic equations for the relevant NS properties. Though it is known that the I-C relation does not hold as good as the I-Love relation, we first focus on it for its simplicity. We further extend our analysis to I-Love relation in later sections and compare the difference of them.
The ordinary differential equations that govern a relativistic, non-rotating star made of perfect fluid are the well-known Tolman-Oppenheimer-Volkov (TOV) equations Oppenheimer and Volkoff (1939); Tolman (1939),
$\frac{{\rm d}r}{{\rm d}p}$ $=$ $-\frac{1}{p+\rho}\frac{r^{2}(1-2m/r)}{m+4\pi r^{3}p}\,$, (1)
$\frac{{\rm d}m}{{\rm d}p}$ $=$ $-\frac{\rho}{p+\rho}\frac{4\pi r^{4}(1-2m/r)}{m+4\pi r^{3}p}\,$, (2)
where $r$ is the circumferential radius and
$m(r)=\int_{0}^{r}4\pi r^{\prime 2}\rho(r^{\prime}){\rm d}r^{\prime}\,$, (3)
can be regarded as the gravitational mass enclosed within radius $r$. Energy density $\rho$ and pressure $p$ are related by the so-called EOS,
$\rho=\rho(p)\,.$ (4)
In the above equation, we have assumed that the density is only a function of the pressure, or to say, we have adopted a cold EOS. The temperature effects are more important for newly born NSs and do affect the universal relations relating the post-merger gravitational wave spectrum with the NS structures Raithel and Most (2022); Han et al. (2025). Nevertheless, for old NSs with stable structures, they are usually cold relative to their Fermi temperature and the above temperature-independent EOS is sufficient.
We calculate the moment of inertia of a NS with the slow rotation approximation Hartle (1967); Hartle and Thorne (1968); Gao et al. (2021). It can be shown that, in this approximation, the moment of inertia of the star also can be described by a single ordinary differential equation Grigorian and Khachatrian (1997) ,
$\frac{{\rm d}i}{{\rm d}p}=-\frac{2}{3}\frac{4\pi r^{6}}{m+4\pi r^{3}p}\left(1-\frac{5}{2}\frac{i}{r^{3}}+\frac{i^{2}}{r^{6}}\right)\,$, (5)
where $i$ can be regarded as the moment of inertial contributed by the matter inside radius $r(p)$. However, due to the relativistic effect, the above equation is nonlinear for $i$ and there is no superposition principle as in Newtonian gravity. There are studies of the universal relations for fast or differential rotating NSs Doneva et al. (2013); Pappas and Apostolatos (2014); Bretz et al. (2015). However, the calculation of moment of inertia in these situations can be much more complex and we leave it for future studies.
With the above equations, the radius, mass and moment of inertia of the star are defined to be
$R=r|_{p=0}\,,\quad M=m|_{p=0}\,,\quad I=i|_{p=0}\,$, (6)
where $p$ vanishes at the boundary of the star.
Figure 1: The I-C relation for NSs and QSs with realistic EOSs as well as polytropic EOSs.
In Fig. 1 we show the I-C universal relation for several realistic EOSs including NSs and QSs as well as polytropic EOSs with $\rho(p)=\rho_{n}(p/\rho_{n})^{n/(n+1)}$. The dimensionless moment of inertial is defined to be $\tilde{I}\equiv I/M^{3}$ and the compactness is $C=M/R$ Yagi and Yunes (2013b, a). As we focus on the dimensionless relation, the dimensionful quantity $\rho_{n}$ dose not affect the final relation and we choose it to be $1$ as the unit. From the figure one can see that for realistic EOSs, the relation holds well to the level of several percentages. However, this relation has different branches for NSs and QSs. As expected, the polytropic EOS with $n=1$ is a good approximation for realistic NS EOSs while $n=0$ can describe the EOSs of QSs.
III Linear response of the I-C relation
In this section, we describe the approach we used to calculate the linear response of arbitrary EOS perturbations. Starting from any given EOS, say, $\rho(p)$, and a given central pressure, $p_{0}$, we may solve the structure equations given in the previous section with the initial condition that
$p=p_{0}\,,\quad r=0\,,\quad m=0\,,\quad i=0\,$, (7)
to get a background solution denoted by
$r(p)\,,\quad m(p)\,,\quad i(p)\,$, (8)
and thus $R$, $M$, II for the star at $p=0$. With the background solution, considering another EOS given by
$\rho^{\prime}(p)=\rho(p)+\delta\rho(p)\,$, (9)
but with the same initial condition, we can write the corresponding NS structure to be
$r^{\prime}(p)$ $=$ $r(p)+\delta r(p)\,$, (10)
$m^{\prime}(p)$ $=$ $m(p)+\delta m(p)\,$, (11)
$i^{\prime}(p)$ $=$ $i(p)+\delta i(p)\,.$ (12)
Assuming $\delta\rho(p)$ is an EOS perturbation satisfying $\delta\rho(p)/\rho(p)\ll 1$ for $p\in[0,p_{0}]$, one can then obtain the linearized differential equations for $\delta r$, $\delta m$ and $\delta i$ by substituting the above equations to the NS structure equations and only keeping the leading order terms. The linearization also requires the variable set $\delta r(p)/r(p)$, $\delta m(p)/m(p)$ and $\delta i(p)/i(p)$ to be small, which is in general guaranteed by the assumption $\delta\rho(p)/\rho(p)\ll 1$. Note that at $p=p_{0}$ the condition $r=m=i=0$ is only a coordinate singularity. At the center of the star, one can solve the TOV equations to obtain
$r(p)=\sqrt{\frac{3}{2\pi}\frac{p_{0}-p}{(\rho_{0}+p_{0})(\rho_{0}+3p_{0})}}+O(p_{0}-p)^{3/2}\,$, (13)
where $\rho_{0}=\rho(p_{0})$. Therefore,
$\left.\frac{\delta r}{r}\right|_{p=p_{0}}=-\frac{\rho_{0}+2p_{0}}{(\rho_{0}+p_{0})(\rho_{0}+3p_{0})}\delta\rho(p_{0})\,$, (14)
which is also small if $\delta\rho(p_{0})/\rho(p_{0})\ll 1$; it is similar for $\delta m/m$ and $\delta i/i$.
The linearized equations for the $\delta$-quantities read
$\frac{{\rm d}}{{\rm d}p}\begin{bmatrix}\delta r\\ \delta m\\ \delta i\end{bmatrix}=\bm{A}(p)\begin{bmatrix}\delta r\\ \delta m\\ \delta i\end{bmatrix}+\bm{S}(p)\delta\rho(p)\,$, (15)
where $\bm{A}(p)$ and $\bm{S}(p)$ are matrix and column vector depending on the background solution. We present the formulae of $\bm{A}(p)$ and $\bm{S}(p)$ here for later numerical calculations,
$\bm{A}(p)$ $=\begin{bmatrix}\dfrac{2}{p+\rho}\dfrac{m^{2}+2\pi r^{4}p-mr-8\pi r^{3}pm}{(m+4\pi r^{3}p)^{2}}&\quad\dfrac{1}{p+\rho}\dfrac{r^{2}(1+8\pi r^{2}p)}{(m+4\pi r^{3}p)^{2}}&\quad 0\\ \dfrac{\rho}{p+\rho}\dfrac{8\pi r^{2}(3m^{2}-2mr-2\pi r^{4}p)}{(m+4\pi r^{3}p)^{2}}&\quad\dfrac{\rho}{p+\rho}\dfrac{4\pi r^{4}(1+8\pi r^{2}p)}{(m+4\pi r^{3}p)^{2}}&\quad 0\\ -\dfrac{4\pi r^{2}(4mr^{3}+8\pi r^{6}p-5im-8\pi pi^{2})}{(m+4\pi r^{3}p)^{2}}&\quad\dfrac{8\pi r^{6}(1-5i/2r^{3}+i^{2}/r^{6})}{3(m+4\pi r^{3}p)^{2}}&\quad\dfrac{4\pi(5r^{3}-4i)}{3(m+4\pi r^{3}p)}\end{bmatrix}\,$, (16)
$\bm{S}(p)$ $=\begin{bmatrix}\dfrac{1}{(p+\rho)^{2}}\dfrac{r^{2}(1-2m/r)}{m+4\pi r^{3}p}\\ -\dfrac{p}{(p+\rho)^{2}}\dfrac{4\pi r^{4}(1-2m/r)}{m+4\pi r^{3}p}\\ 0\end{bmatrix}\,$. (17)
The initial condition for the perturbation equation, Eq. (15), reads,
$p=p_{0}\,,\quad\delta r=0\,,\quad\delta m=0\,,\quad\delta i=0\,.$ (18)
A further observation leads to some intuitive simplification of the problem. Note that the perturbation equation (15) is linear, thus a natural and complete choice of $\delta\rho(p)$ is a delta function that $\delta\rho(p)=\delta(p-p_{1})$ with $0<p_{1}\leq p_{0}$. Considering this kind of EOS perturbation, Eq. (15) is simplified into
$\frac{{\rm d}}{{\rm d}p}\begin{bmatrix}\delta r\\ \delta m\\ \delta i\end{bmatrix}=\bm{A}(p)\begin{bmatrix}\delta r\\ \delta m\\ \delta i\end{bmatrix}\,$, (19)
but with initial condition
$p=p_{1}\,,\quad\delta r=-S_{1}(p_{1})\,,\quad\delta m=-S_{2}(p_{1})\,,\quad\delta i=0\,$, (20)
where we have defined $\bm{S}(p)=\big[S_{1}(p),S_{2}(p),0\big]^{\intercal}$.
Before further proceeding, here we note that the dimension of the delta function source $\delta\rho=\delta(p-p_{1})$ is $p^{-1}$ instead of $\rho$ according to the definition of delta function. Therefore, the corresponding solution like $\delta r$ has a dimension of $r/p\rho$ instead of $r$, which might be less intuitive. Nevertheless, physical quantities that are obtained by applying superposition principle have the correct dimension as shown below.
Solving the above equation analytically or numerically, one can obtain the perturbation quantities as a function of $p$. We denote the solution as
$\delta r(p;p_{1})\,,\quad\delta m(p;p_{1})\,,\quad\delta i(p;p_{1})\,$, (21)
where $p_{1}$ reminds that the solution is for $\delta\rho(p)=\delta(p-p_{1})$.
What we are interested in is the change in the star’s radius, mass and moment of inertial caused by an arbitrary EOS perturbation. As we have linearized the equations, we can apply the superposition principle to calculate; for example,
$\delta R=\int_{0}^{p_{0}}\delta R(p_{1})\delta\rho(p_{1}){\rm d}p_{1}\,$, (22)
where $\delta R(p_{1})\equiv\delta r(p=0;p_{1})$. The deviation of the universal relation then can be characterized by the following combination,
$\frac{{\rm D}\tilde{I}}{\tilde{I}}\equiv\frac{\delta\tilde{I}-K\delta C}{\tilde{I}}\,$, (23)
where
$\delta\tilde{I}/\tilde{I}=\delta I/I-3\delta M/M^{3}\,,\quad\delta C/C=\delta M/M-\delta R/R\,$, (24)
and $K\equiv{\rm d}\tilde{I}/{\rm d}C$ is the “true” slope of the universal solution. Notice that it is similar to Eq. (6.30) in Ref. Chan et al. (2016) but we do not expand to the leading order of $C$. However, due to the fact that the relation is only approximately universal, $K$ can be calculated with the background EOS.
Equation (23) is a linear combination of the perturbation quantities, thus it also adopts a superposition form
$\frac{{\rm D}\tilde{I}}{\tilde{I}}=\int_{0}^{1}p_{0}\rho_{0}\frac{{\rm D}\tilde{I}(p_{1})}{\tilde{I}}\frac{\delta\rho(p_{1})}{\rho_{0}}{\rm d}x\,$, (25)
where $x=p_{1}/p_{0}$, and ${\rm D}\tilde{I}(p_{1})/\tilde{I}$ is the corresponding quantity for a delta function like EOS perturbation defined before. $\rho_{0}$ can be any number that makes these combination dimensionless, and we choose it to be $\rho_{0}=\rho(p_{0})$ as it seems to be the most natural. The dimensionless factor $p_{0}\rho_{0}{\rm D}\tilde{I}(p_{1})/\tilde{I}$ then fully characterizes the linear response of the I-C relation to an arbitrary EOS perturbation at a given compactness. As a function of $p_{1}$, it will also tell us which part of the NS is most sensitive to the EOS perturbation and thus contributes the most.
Figure 2: Illustration of the equivalence between the EOS perturbation and central pressure perturbation for a truly universal relation.
It is interesting and intuitive to consider another perturbation here: the perturbation of the central pressure. If the relation we consider is truly universal, one can conclude that any EOS perturbation should be equivalent to some central pressure perturbation with an unperturbed EOS, in the sense that they all give points in the unique curve and there is a one to one correspondence between the central pressure and the point in this curve. We call this as local universality and give an illustration in Fig. 2.
We will consider another star that has EOS $\rho=\rho(p)$ but with the initial condition
$p=p_{0}+\Delta p\,,\quad r^{\prime\prime}=0\,,\quad m^{\prime\prime}=0\,,\quad i^{\prime\prime}=0\,.$ (26)
Here we use $\Delta$” to distinguish it from the perturbation of EOS. To deal with this perturbation, we can integrate one step outwards to get
$p=p_{0}\,,\quad r^{\prime\prime}=\Delta r_{0}\propto|\Delta p|^{1/2}\,,\quad m^{\prime\prime}=0\,,\quad i^{\prime\prime}=0\,$, (27)
where we have only kept the leading terms, so that $m^{\prime\prime}\propto|\Delta p|^{3/2}$ and $i^{\prime\prime}\propto|\Delta p|^{5/2}$ are ignored. Taking this as the initial condition, we can write the corresponding solution as
$r^{\prime\prime}(p)$ $=$ $r(p)+\Delta r(p)\,$, (28)
$m^{\prime\prime}(p)$ $=$ $m(p)+\Delta m(p)\,$, (29)
$i^{\prime\prime}(p)$ $=$ $i(p)+\Delta i(p)\,.$ (30)
And again one can derive the equation for the perturbation
$\frac{{\rm d}}{{\rm d}p}\begin{bmatrix}\Delta r\\ \Delta m\\ \Delta i\end{bmatrix}=\bm{A}(p)\begin{bmatrix}\Delta r\\ \Delta m\\ \Delta i\end{bmatrix}\,$. (31)
Note that Eq. (31) and Eq. (19) are exactly the same equation, but with different initial conditions. Further, as they are linear equations, scaling the initial condition simply scales the final solution. We can rescale the initial condition for Eq. (19) to be
$p=p_{1}\,,\quad\delta r=-r_{1}\,,\quad\delta m=4\pi r_{1}^{3}p_{1}\,,\quad\delta i=0\,$, (32)
with $r_{1}$ the radius where $p=p_{1}$. Note that when $p_{1}\rightarrow p_{0}$, this initial condition is just the initial condition in Eq. (27). Thus, we can say that the EOS perturbation at the center of the star is equivalent to the central pressure perturbation.
The above discussion gives us another method of calculating $K$ in Eq. (23),
$K\equiv\frac{{\rm d}\tilde{I}}{{\rm d}C}=\frac{\Delta\tilde{I}}{\Delta C}=\frac{\delta\tilde{I}(p_{1}=p_{0})}{\delta C(p_{1}=p_{0})}\,$, (33)
which means that the factor $p_{0}\rho_{0}{\rm D}\tilde{I}(p_{1})/\tilde{I}$ is always zero at the center of the star. A necessary condition for the relation to be truly universal is $\delta\tilde{I}(p_{1})/\delta C(p_{1})=\Delta\tilde{I}/\Delta C$ irrespective of $p_{1}$, so that the factor $p_{0}\rho_{0}{\rm D}\tilde{I}(p_{1})/\tilde{I}$ is always zero throughout the star.
IV Analytical solution to a constant-density star in the Newtonian gravity
It is suggest that the universal relations may come from the similarity of the realistic EOSs and an incompressible, constant-density EOS Sham et al. (2015). Previous study also showed that, the I-Love relation is independent of leading order EOS perturbation around the incompressible EOS in the Newtonian limit (at the relative leading order in CCChan et al. (2016). Thus here we study the I-C relation for a constant-density star in the Newtonian gravity analytically as an illustration.
In the Newtonian gravity, after taking a dimensionless form for convenience, Eq. (19) is simplified into
$\frac{{\rm d}}{{\rm d}p}\begin{bmatrix}\dfrac{\delta r}{r}\\ \dfrac{\delta m}{m}\\ \dfrac{\delta i}{i}\end{bmatrix}=\begin{bmatrix}-\dfrac{r}{\rho m}&\dfrac{r}{\rho m}&0\\ -\dfrac{16\pi r^{4}}{m^{2}}&\dfrac{8\pi r^{4}}{m^{2}}&0\\ -\dfrac{16\pi r^{6}}{im}&\dfrac{8\pi r^{6}}{3im}&\dfrac{8\pi r^{6}}{3im}\end{bmatrix}\begin{bmatrix}\dfrac{\delta r}{r}\\ \dfrac{\delta m}{m}\\ \dfrac{\delta i}{i}\end{bmatrix}\,$, (34)
with initial conditions
$p=p_{1}\,,\quad\delta r/r=-r_{1}/m_{1}\rho_{1}^{2}\,,\quad\delta m/m=0\,,\quad\delta i/i=0\,$, (35)
where $r_{1}=r(p_{1})$ and $m_{1}=m(p_{1})$. For a constant-density star, one has
$4\pi r^{3}\rho/m=3\,,\quad 8\pi r^{5}\rho/i=15\,.$ (36)
Thus, by introducing the following combination,
$\delta u$ $=$ $\frac{\delta r}{r}-\frac{1}{4}\frac{\delta m}{m}\,$, (37)
$\delta v$ $=$ $\frac{\delta r}{r}-\frac{1}{3}\frac{\delta m}{m}\,$, (38)
$\delta w$ $=$ $\frac{\delta i}{i}-20\delta u+15\delta v\,$, (39)
Eq. (34) can be directly solved as
$\delta u(p;p_{1})$ $=$ $-\frac{r_{1}}{m_{1}\rho_{1}^{2}}e^{\int_{p_{1}}^{p}\frac{2r}{\rho m}{\rm d}p}\,$, (40a)
$\delta v(p;p_{1})$ $=$ $-\frac{r_{1}}{m_{1}\rho_{1}^{2}}e^{\int_{p_{1}}^{p}\frac{3r}{\rho m}{\rm d}p}\,$, (40b)
$\delta w(p;p_{1})$ $=$ $\frac{5r_{1}}{m_{1}\rho_{1}^{2}}e^{\int_{p_{1}}^{p}\frac{5r}{\rho m}{\rm d}p}\,.$ (40c)
Note that the integral
$\int_{p_{1}}^{p}\frac{r}{\rho m}{\rm d}p=\int_{r}^{r_{1}}\frac{{\rm d}r}{r}=\ln\frac{r_{1}}{r}\,$, (41)
for a constant-density star. Combining with Eq. (40), we have
$\frac{{\rm D}\tilde{I}(p_{1})}{\tilde{I}}$ $=$ $\frac{\delta I(p_{1})}{I}-3\frac{\delta M(p_{1})}{M}-K\frac{C}{\tilde{I}}\left(\frac{\delta M(p_{1})}{M}-\frac{\delta R(p_{1})}{R}\right)$ (42)
$=$ $\delta w(p=0;p_{1})+3\delta v(p=0;p_{1})$
$=$ $-\frac{3r_{1}}{m_{1}\rho^{2}}\left(\frac{r_{1}}{R}\right)^{3}+\frac{5r_{1}}{m_{1}\rho^{2}}\left(\frac{r_{1}}{R}\right)^{5}\,$,
where we have used $K=-2\tilde{I}/C$ for a constant-density star in the Newtonian gravity. Taking into account that
$m(r)$ $=$ $\frac{4\pi}{3}r^{3}\rho\,$, (43)
$r(p)$ $=$ $\sqrt{\frac{3}{2\pi\rho^{2}}}(p_{0}-p)^{1/2}\,$, (44)
we have
$p_{0}\rho_{0}\frac{{\rm D}\tilde{I}(p_{1})}{\tilde{I}}=\left(1-\frac{p_{1}}{p_{0}}\right)^{1/2}\left(1-\frac{5}{2}\frac{p_{1}}{p_{0}}\right)\,.$ (45)
Interestingly, the above factor only depends on the ratio $p_{1}/p_{0}$ thus is universal for all background stars. As we will see later, for realistic EOSs, this factor in general also depends on the compactness of the background star. Though it is in the Newtonian gravity and for a constant-density star, Eq. (45) still captures the order of magnitude of this factor, thus provides a very intuitive estimation of the universality of the I-C relation.
V Linear response for compact stars with realistic EOSs
In this section, we focus on NSs and QSs with realistic EOSs in GR. We use the EOS data provided by Lattimer and Prakash (2001), which are in a tabulated form. The EOSs for NS have a crust EOS down to a density of $10^{8}\,{\rm g\,cm^{-3}}$. To obtain the factor $p_{0}\rho_{0}{\rm D}\tilde{I}(p_{1})/\tilde{I}$ for realistic EOSs, in general one needs to numerically solve Eq. (19). In Fig. 3, we show this factor for NSs with the AP4 EOS Akmal et al. (1998) as well as QS with the SQM1 EOS Prakash et al. (1995). As in general this factor depends on the background stars, we plot the NS cases for three different star compactnesses. We also plot Eq. (45) for comparison. One can see that, the Newtonian constant density approximation does capture the order of magnitude for the factor in realistic cases, except for a small region where $p_{1}/p_{0}$ is close to zero, that is, the outer region of the NS. We find that when $p_{1}/p_{0}$ tends to zero, this factor grows up towards infinity (or towards a relatively large but finite number in the case of QSs) for realistic EOSs. We will discuss this behavior later in detail as it is closely related to the fact that NSs and QSs have different branches of the I-C relation.
Figure 3: The factor $p_{0}\rho_{0}{\rm D}\tilde{I}/\tilde{I}$ as a function of $p_{1}/p_{0}$. Note that unlike the analytical case we discussed before, for realistic EOSs, the factor in general depends on the background stars. Here we show the factor for NSs with the AP4 EOS and different compactnesses as well as for QSs with the SQM1 EOS. We also show this factor for a polytropic EOS with $n=0$ in the Newtonian gravity given in Eq. (45).
Figure 4: The factor $\delta\rho/\rho_{0}$ for different realistic EOSs relative to the AP4 EOS. The background star has a compactness $C=0.2$. Different from the NS-NS case, the difference between the QS EOS and the NS EOS is finite at $p_{1}/p_{0}=0$.
Figure 5: The product of the two factors from Fig. 3 and Fig. 4. The relative deviation then can be “read off” from the figure as the area enclosed by the curve and the $x$-axis.
Figure 6: The integration cumulated from the center of the star to its surface.
To estimate the integral in Eq. (25), which represents the deviation of the relation caused by the difference between realistic EOSs, in Fig. 4 and Fig. 5, we show the factor $\delta\rho(x)/\rho_{0}$ and the product of the two factors for some realistic EOS perturbations, respectively. Here $\delta\rho$ represents the difference between the AP4 EOS and the other EOSs. From the figures, one can easily estimate that the universal relation will hold at a percent level. The sign change further minishes the integral. One can also roughly conclude that the mean contribution of the violation of the relation comes from the outer region of the star. This is shown in Fig. 6, where we plot the integration cumulated from the center of the star where $p_{1}/p_{0}=1$ to its surface where $p_{1}/p_{0}=0$. It is clear that, when considering the deviation caused by the EOS difference between a QS EOS and a background NS EOS, the integral becomes divergent at the surface of the star, which is caused by the finite density difference at $p=0$ for a NS and a QS as well as the divergent behavior of the factor $p_{0}\rho_{0}{\rm D}\tilde{I}(p_{1})/\tilde{I}$.
Figure 7: The behavior of different factors that compose ${\rm D}\tilde{I}/\tilde{I}$ for the AP4 EOS and a star compactness $C=0.2$. $f$ in the $y$-axis label represents for $\delta R/R$, $\delta M/M$ and so on. The divergent behavior comes from $\delta R/R$ and can be understood with the initial condition $\delta r/r$.
To understand this divergent behavior, we notice that in Eq. (42), the behavior of the factor $p_{0}\rho_{0}{\rm D}\tilde{I}(p_{1})/\tilde{I}$ depends on $\delta I(p_{1})/I$, $\delta M(p_{1})/M$, and $\delta R(p_{1})/R$, which are directly solved from the perturbation equations. In Fig. 7, we plot these factors separately. From this figure, one can clear see that the divergent behavior at the surface of the star is dominated by the factor $\delta R(p_{1})/R$.
The divergence of $\delta R(p_{1})$ near the surface of the NS can be analytically estimated via a polytropic crust EOS model. Considering EOS perturbations in the low pressure region, the inner structure of the star then is irrelevant to our problem and the effect can be fully characterized by the statement that, at $p=p_{R}$ and $\rho=\rho_{R}$, the core of the star has mass $M$, radius $R$, where $p_{R}$ is some small pressure. We can then assume that for $p<p_{R}$, the NS crust has a polytropic EOS that $\rho(p)=\rho_{R}(p/p_{R})^{n/(n+1)}$ and we can model the crust of the NS with Newtonian approximation,
$\frac{{\rm d}r}{{\rm d}p}=-\frac{1}{\rho}\frac{r^{2}}{m}\,.$ (46)
Further, it is known that the crust of NS contains only a small amount of mass but provides a relatively considerable radius Chan et al. (2015), therefore, when considering the behavior of $\delta r$, we can set $m=M$ in the above equation (also note that $\delta r$ is the only divergent part so we can safely ignore $\delta m$ there).
With these approximation, one can solve the structure of the curst analytically, which reads
$\frac{1}{r}-\frac{1}{R}=(n+1)\frac{p_{R}^{\frac{n}{n+1}}}{M\rho_{R}}\left(p^{\frac{1}{n+1}}-p_{R}^{\frac{1}{n+1}}\right)\,.$ (47)
This gives a crust with a depth $z\approx(n+1)R^{2}p_{R}/M\rho_{R}$. Following the procedure introduced in Sec. III, one can then write down the equation for $\delta r(p)$ caused by an EOS perturbation $\delta\rho(p)=\delta(p-p_{1})$ as
$\frac{{\rm d}\delta r}{{\rm d}p}=-\frac{2r}{\rho M}\delta r\,$, (48)
with initial condition $\delta r(p_{1})=-r(p_{1})^{2}/\rho(p_{1})^{2}M$. This equation again can be integrated analytically and one can obtain
$\delta R(p_{1})=\delta r(p_{1})\left(\frac{\frac{M\rho_{R}/R-p_{R}}{p_{R}^{n/(n+1)}}+p_{1}^{1/(n+1)}}{\frac{M\rho_{R}/R-p_{R}}{p_{R}^{n/(n+1)}}}\right)^{2}\,\approx-\frac{R^{2}}{M\rho(p_{1})^{2}}.$ (49)
Thus, near the surface $\delta R(p_{1})/R$ diverges as $\rho(p_{1})^{-2}$. In Fig. 7, we also plot the above estimation with $M/R\approx 0.2$. It can be seen that this estimation captures the divergent behavior of $\delta R(p_{1})/R$. Differently for QSs, near the surface, they behave more like the constant density star as we discussed before. Factors remain finite there.
We should note that, the divergence of $\delta R(p_{1})$ is a physical result and itself does not cause problem to our linear analysis. The quantity that needs to be finite and small is the integral defined in Eq. (22) when the linear approximation holds. As an example, we may consider the EOS perturbation with $n\rightarrow n+\epsilon$ in the above crust model. The corresponding change in the density is
$\delta\rho(p)=\frac{\epsilon\rho_{R}}{(n+1)^{2}}\left(\frac{p}{p_{R}}\right)^{\frac{n}{n+1}}\ln\frac{p}{p_{R}}\,$, (50)
which is well defined for $n>0$. The integration,
$\int_{0}^{p_{R}}{\rm d}p_{1}\delta R(p_{1})\delta\rho(p_{1})\approx\epsilon\frac{R^{2}p_{R}}{M\rho_{R}}\,$, (51)
is just the change in $z$ that $\delta z=\delta nR^{2}p_{R}/M\rho_{R}=\epsilon R^{2}p_{R}/M\rho_{R}$. One can also notice that, in Fig. 6, the integration of ${\rm D}\tilde{I}$ is finite for NS-NS comparison. As $\delta R(p_{1})$ is the only divergent part in ${\rm D}\tilde{I}$, the integration of $\delta R(p_{1})$ for realistic NS EOSs, remains finite.
Figure 8: The behavior of different factors that compose ${\rm D}\tilde{I}/\tilde{I}$ for the SQM1 EOS and a star compactness $C=0.2$.
Finally, the integrated factor $p_{0}\rho_{0}{\rm D}\tilde{I}/\tilde{I}$ is divergent if one considers a NS background but takes a QS EOS as the perturbed EOS. As the QS EOS has a finite surface density $\rho_{s}$ while NS EOS at the low density part can be approximated by $\rho\propto p^{3/5}$, the integration at small $p_{1}$ is roughly $\int 1/\rho_{1}^{2}{\rm d}p_{1}$ and divergent. As we will see in Sec. VII, this divergence means that the linear approximation fails for this situation, suggesting an intrinsic difference in the I-C relation for NSs and QSs. We also note that, the equivalence between taking any one of the star as background only holds when both choices give finite results and it does not apply to the NS-QS comparison.
VI I-Love relation
It is known that the I-Love universal relation holds better compared to the I-C relation. Previous study showed that the leading order dependence on $C$ cancels out for perturbation of the I-Love relation Chan et al. (2016). In this section, we apply our approach introduced in previous sections to the I-Love relation.
Tidal love number in GR can be calculated following the procedure developed by Hinderer (2008), and here we list the relevant results for convenience. The tidal deformability is,
$\Lambda=\frac{2}{3}k_{2}R^{5}\,$, (52)
where
$k_{2}$ $=$ $\frac{8C^{5}}{5}(1-2C)^{2}\big[2+2C(Y-1)-Y\big]$ (53)
$\times$ $\Big\{2C\big[6-3Y+3C(5Y-8)\big]$
$+$ $4C^{3}\big[13-11Y+C(3Y-2)+2C^{2}(1+Y)\big]$
$+$ $3(1-2C)^{2}\big[2-Y+2C(Y-1)\big]\ln(1-2C)\Big\}^{-1}\,$,
is the dimensionless Love number; $Y=y\big|_{r=R}$ needs to be calculated from the ordinary differential equation
$\frac{{\rm d}y}{{\rm d}r}$ $=$ $\left(\frac{1}{r}-A\right)y-\frac{y^{2}}{r}-rB\,$, (54)
where
$A$ $=$ $\frac{2}{r}+\frac{1}{1-2m/r}\left[\frac{2m}{r}+4\pi r(p-\rho)\right]\,$, (55)
$B$ $=$ $-\frac{6}{r^{2}(1-2m/r)}+\frac{4\pi}{1-2m/r}\Bigg[5\rho+9p$ (56)
$+$ $(\rho+p)\frac{{\rm d}\rho}{{\rm d}p}\Bigg]-\left(2\frac{m+4\pi r^{3}p}{r^{2}(1-2m/r)}\right)^{2}\,$,
with an initial value $y\big|_{r=0}=2$. In Fig. 9 we show the I-Love relation for several realistic EOSs as well as polytropic EOSs. This universal relation holds better compared to the I-C relation at $\sim 0.1\%$ level. This relation is also universal among NSs and QSs.
Figure 9: The I-Love relation for NSs and QSs with realistic EOSs as well as polytropic EOSs. The dimensionless moment of inertial is defined to be $\tilde{I}\equiv I/M^{3}$ and the dimensionless tidal deformability $\tilde{\Lambda}\equiv\Lambda/M^{5}$.
By adding an EOS perturbation with a form $\delta\rho=\delta(p-p_{1})$ and solving the perturbation quantities to the leading order, one can obtain a similar factor $\delta\Lambda/\Lambda$. The concerned quantity now is
$\left(\frac{{D}\tilde{I}}{\tilde{I}}\right)_{\tilde{\Lambda}}\equiv\frac{\delta\tilde{I}-K_{\tilde{\Lambda}}\delta\tilde{\Lambda}}{\tilde{I}}\,$, (57)
where we use the subindex $\tilde{\Lambda}$ to distinguish it from the quantity before. However, due to the term ${\rm d}\rho/{\rm dp}$ in the differential equation, additional care needs to be taken for calculating $\delta\Lambda/\Lambda$. In Appendix A, we give the detailed calculation procedure for this factor.
Though for the general cases in GR, the numerical calculation takes some efforts, it is still easy and intuitive to have an analytical solution for this factor in the Newtonian gravity for a constant-density star as we do for the I-C relation. Remind that in the Newtonian gravity, the tidal Love number is governed by the Clairaut-Radau equation Poisson and Will (2014),
$r\frac{{\rm d}\eta_{l}}{{\rm d}r}+\eta_{l}(\eta_{l}-1)+6\frac{4\pi r^{3}\rho}{3m}(\eta_{l}+1)-l(l+1)=0\,$, (58)
with an initial condition $\eta_{l}\big|_{r=0}=l-2$. The dimensionless Love number is given by
$k_{l}=\frac{l+1-\eta_{l}(R)}{2\big[l+\eta_{l}(R)\big]}\,.$ (59)
For a constant-density star and the $l=2$ Love number that we are interested, the background solution is simply $\eta_{2}(r)=0$ and $k_{2}=3/4$. Perturbing Eq. (58) with $\delta\rho(p)=\delta(p-p_{1})$ and simplifying with the background solution, one gets
$\frac{{d}\delta\eta}{{\rm d}p}=\frac{18r}{\rho m}\frac{{\rm d}r}{r}-\frac{6r}{\rho m}\frac{\delta m}{m}+\frac{5r}{\rho m}\delta\eta\,$, (60)
with initial conditions
$p=p_{1}\,,\quad\delta\eta=-\frac{6r_{1}}{m_{1}\rho_{1}^{2}}\,.$ (61)
The combination
$\delta z=\delta\eta+9\frac{\delta r}{r}-3\frac{\delta m}{m}\,$, (62)
then solves the equation with
$\delta z(p;p_{1})=-\frac{15r_{1}}{m_{1}\rho_{1}^{2}}e^{\int_{p_{1}}^{p}\frac{5r}{\rho m}{\rm d}p}\,.$ (63)
Using the fact that
$\frac{\delta\Lambda}{\Lambda}=\frac{\delta k_{2}}{k_{2}}+5\frac{\delta R}{R}=-\frac{5}{6}\delta\eta+5\frac{\delta R}{R}\,$, (64)
and $K_{\tilde{\Lambda}}=2/5$ for a constant-density star in the Newtonian gravity and taking into account the solutions in Eq. (40), one can easily verify that, the factor defined in Eq. (57) is
$\left(\frac{{\rm D}\tilde{I}(p_{1})}{\tilde{I}}\right)_{\tilde{\Lambda}}=0\,.$ (65)
This result is in fact equivalent to the result obtained by Chan et al. (2016), where they expanded the EOS with $\rho(p)=c_{0}+c_{1}p+c_{2}p^{2}+\cdots$ and expanded the star structure with $C$. They found that the I-Love relation does not depend linearly in $c_{n}$ to the leading order of $C$ for $n\neq 0$. As the polynomial expansion of EOSs does span all the possible perturbations and the leading order of $C$ is equivalent to the Newtonian limit, while linear in $c_{n}$ is linear in the EOS perturbation, their result has exactly the same meaning as we obtain here. However, we get this result with a different approach that is more convenient to consider arbitrary realistic EOS perturbations and can be simply extended to full GR cases.
Figure 10: The factor $p_{0}\rho_{0}({\rm D}\tilde{I}/\tilde{I})_{\tilde{\Lambda}}$ as a function of $p_{1}/p_{0}$. As before, we plot this factor for the AP4 and SQM1 EOSs. Here we also plot the factor for the constant-density star in GR with $C=0.2$, while its Newtonian counterpart leads to zero. This factor is similar to the factor for the I-C relation but is an order of magnitude smaller, providing a better universality in the I-Love relation.
Figure 10 shows the factor $p_{0}\rho_{0}({\rm D}\tilde{I}/\tilde{I})_{\tilde{\Lambda}}$ for several different background stars in GR. It can be seen that, compared to the factor for the I-C relation, this factor is about an order of magnitude smaller, thus provides a much better universality. Different from the factor for the I-C relation, here for QSs, at the surface of the star the factor seems approaching zero, leading to a weaker response to the EOS perturbation, which is consistent with the suggestion that the I-Love relation comes from the constant-density property Sham et al. (2015). We also find that, for a polytropic EOS with $n=0$ in GR and $C=0.2$, this factor is relatively small but not strictly zero. However, this smallness may not be able to be simply attributed as a continuation from the Newtonian limit. $C=0.2$ is already a large compactness if one notices that, the second order term in $C$ for the expansion of $\tilde{\Lambda}$ given in Ref. Chan et al. (2016), Eq. (6.28), is already larger than the leading order term in this case.
VII Validity of the linear order analysis
Figure 11: The linear perturbation prediction for the I-C relation for EOS perturbations that continuously connect the AP4 EOS and the MPA1 EOS. The black line denotes the linear perturbation prediction while the red dashed line is the tangent line at the background point. The circles mark the star corresponding to the same central pressure as the background star for each EOS, while the triangles mark the corresponding linear prediction.
Figure 12: Similar to Fig. 11, but for the I-Love relation.
In previous sections, we discuss the linear response of the I-C and I-Love universal relations under EOS perturbations. One natural question is that to what extend this linear approximation holds and whether it can be applied to realistic EOSs. In this section, we discuss this question based on examples of realistic EOS differences.
Taking AP4 and MPA1 as example, in Fig. 11 and Fig. 12, we plot the linear perturbation prediction for these two EOSs. Defining
$\delta\rho(p)=\rho_{\rm MPA1}(p)-\rho_{\rm AP4}(p)\,$, (66)
we further construct a series of 11 EOSs that connect the AP4 EOS and the MPA1 EOS via
$\rho_{i}(p)=\rho_{\rm AP4}(p)+i\delta\rho(p)/10\,,\quad i=0,\cdots,10\,.$ (67)
In Fig. 11 and Fig. 12, we plot the relation for all these EOSs. Taking the star with a compactness $C=0.2$ and the AP4 EOS as the background star, we also denote the stars with the same central pressure in different EOSs. We can see that, the linear theory predicts the perturbation well for small EOS deviations, however, the real difference between the AP4 EOS and the MPA1 EOS seems to be out of the linear range. We also note that, if the relation is truly universal, the black line should coincide with the red dashed line. As expected, the I-Love relation shows better universality at the linear order.
Figure 13: Similar to Fig. 11, but for the SQM1 EOS. To show the deviation behavior clearly, we plot three groups of lines as introduced in the main text.
We also show an example for NS-QS in Fig. 13 and Fig. 14, in which we take the EOS perturbation as the difference between the SQM1 EOS and the AP4 EOS. For the I-C relation, as discussed before, from our linear theory the deviation caused by such an EOS perturbation will diverge due to the small $p$ behavior in the factor (23). This divergence is caused by the factor $\delta R(p_{1})/R$ and in fact leads to a divergent estimation of $\delta C$. It is also clearly shown in Fig. 13, which is similar to Fig. 11 but where we plot three groups of lines for clarity. They correspond to EOSs that connect the AP4 EOS and the SQM1 EOS via
$\rho_{i,k}(p)$ $=$ $\rho_{\rm AP4}(p)+i\delta\rho(p)/10^{k}\,$, (68)
$i=0,\cdots,10\,,\quad k=1,2,3\,$,
where $\delta\rho=\rho_{\rm SQM1}-\rho_{\rm AP4}$. We can see that, when the EOS deviates from AP4 to SQM1, the change in the I-C relation does not depend on the EOS perturbation linearly, which should behave like in Fig. 13. The linear theory predicts this nonlinear nature, as shown by the black line corresponding to a divergent $\delta C$, but cannot predict the correct finite deviation that is dominated by nonlinear effects in this case. Starting from the QS EOS might be a more natural choice but also cannot connect to the NS EOS. As we discussed before, the QS EOS has a finite density at the surface of the star so that the integral will not diverge. The figure also shows that close to the QS, the deviation caused by the EOS perturbation behaves relatively normally.
Figure 14: Similar to Fig. 12, but for the SQM1 EOS.
The above figures all show that our linear analysis seems unable to capture the difference between realistic EOSs as their difference is too large. However, one aspect we have not used before is that, when considering the universal relations among dimensionless quantities, we have a scaling freedom that corresponds to changing the remaining unit of length in the geometric units. Namely, an EOS with
$\rho^{\prime}(p)=\alpha\rho(p/\alpha)\,$, (69)
will provide exactly the same relation as the EOS $\rho(p)$, where $\alpha$ is any positive constant. We have already encountered this fact for polytropic EOSs as we choose $\rho_{n}=1$. Therefore, a proper scaling may be applied before the linear analysis as the scaling changes the meaning of the same central pressure. In fact, for I-Love relation, a comparison should be made between the dimensionless moment of inertia of two stars with different EOSs but the same dimensionless tidal deformability. The quantities ${\rm D}\tilde{I}$ we defined in Eq. (57), originating from the rule of partial differentiation
$\left.\frac{\partial\tilde{I}(\tilde{\Lambda}(p_{c},\mathrm{EOS}),\mathrm{EOS})}{\partial\mathrm{EOS}}\right|_{\tilde{\Lambda}}=\left.\frac{\partial\tilde{I}}{\partial{\mathrm{EOS}}}\right|_{p_{c}}-\left.\frac{\partial\tilde{I}}{\partial\tilde{\Lambda}}\right|_{\mathrm{EOS}}\left.\frac{\partial\tilde{\Lambda}}{\partial\mathrm{EOS}}\right|_{p_{c}}\,$, (70)
only provide such a comparison at the linear order. It is possible that realistic EOSs that largely deviate from each other might be similar after scaling. Therefore, for comparing realistic EOSs, we suggest that a proper scaling needs to be done before applying the linear analysis. From previous figures we find that, the linear prediction of $\delta\tilde{I}$ is in fact pretty accurate even for realistic EOS differences. Thus, a scaling that leads to the comparison of two stars with a similar $C$ or $\tilde{\Lambda}$ should be well described by the linear response.
VIII Conclusion
In the present paper we studied the linear response of I-C and I-Love universal relations under arbitrary EOS perturbations. We separated the deviation of the universal relations caused by an EOS perturbation into a product of $\delta\rho/\rho_{0}$ and $p_{0}\rho_{0}{\rm D}\tilde{I}/\tilde{I}$, where the second factor only depends on the background star configurations so that it provides an overall estimation of the universality. We analytically calculated this factor for a constant-density star in the Newtonian gravity, and our result for the I-Love relation is consistent with the result in Ref. Chan et al. (2016). We further applied our procedure to realistic EOSs in GR and the linear response analysis estimates the order of magnitude of universality correctly. We found that, for the I-C relation, the divergent behavior when comparing the QS EOS and the NS EOS originates from the radius response $\delta R/R$, and this fully nonlinear behavior suggests that NSs and QSs follow different I-C relations, consistent with real scenarios. We also discussed the validity of applying the linear analysis to study realistic EOS differences and we suggested that a proper scaling of EOSs should be made before applying the procedure.
Our study can be regarded as a new framework for quantitative representation of the universality for NSs. The analysis in this work may not be able to point out the possible origin of the universality but gives an quantitative and intuitive way to describe how good the universality is. Analytical results we obtained in the Newtonian gravity thus can serve as a rough estimation while our analysis can be applied in full GR cases. In this work we only focused on the I-C and I-Love universal relations. The extension to, for example, the I-Love-Q relation, can be done without conceptual difficulties. Including any new quantity basically adds one new perturbation equation that will depend on, but not influence, the other equations as we do for $\Lambda$. We hope that our work reported here could trigger other investigations along this direction.
Acknowledgements.
We thank Kent Yagi and Takuya Katagiri for helpful discussions, and the anonymous referee for useful comments. This work was supported by the National SKA Program of China (2020SKA0120300), the National Natural Science Foundation of China (124B2056, 12573042), the Beijing Natural Science Foundation (1242018), the Max Planck Partner Group Program funded by the Max Planck Society, and the High-Performance Computing Platform of Peking University. Z.H. is supported by the China Scholarship Council (CSC).
Appendix A Perturbative calculation of $\delta\Lambda/\Lambda$
In this appendix, we give a relatively detailed calculation procedure for the linear response of $\Lambda$ caused by EOS perturbations. Starting from the differential equation for $y$,
$\frac{{\rm d}y}{{\rm d}p}=f\left(r,m,y,\rho,\frac{{\rm d}\rho}{{\rm d}p}\right)\,$, (71)
where we have changed the differential variable from $r$ to $p$ and have used $f$ to denote the complex right hand side of the equation. We explicitly write out the dependence on ${\rm d}\rho/{\rm d}p$ in $f$ as well. Consider any EOS perturbation $\rho^{\prime}(p)=\rho(p)+\delta\rho(p)$, we can expand Eq. (71) for a given background solution. Denoting
$y^{\prime}(p)=y(p)+\delta y(p)\,$, (72)
at the leading order, the equation for $\delta y$ reads
$\frac{{\rm d}\delta y}{{\rm d}p}$ $=$ $\frac{\partial f}{\partial r}\delta r+\frac{\partial f}{\partial m}\delta m+\frac{\partial f}{\partial y}\delta y$ (73)
$+\frac{\partial f}{\partial\rho}\delta\rho+\frac{\partial f}{\partial({\rm d}\rho/{\rm d}p)}\frac{{\rm d}\delta\rho}{{\rm d}p}\,.$
This is a linear equation in $\delta$-quantities, however, considering $\delta\rho=\delta(p-p_{1})$, we further rewrite the equation as
$\frac{{\rm d}\delta\tilde{y}}{{\rm d}p}$ $=$ $\frac{\partial f}{\partial r}\delta r+\frac{\partial f}{\partial m}\delta m+\frac{\partial f}{\partial y}\delta\tilde{y}$ (74)
$+\left(\frac{\partial f}{\partial\rho}-\frac{{\rm d}}{{\rm d}p}\frac{\partial f}{\partial({\rm d}\rho/{\rm d}p)}+\frac{\partial f}{\partial y}\frac{\partial f}{\partial({\rm d}\rho/{\rm d}p)}\right)\delta\rho\,$,
where
$\delta\tilde{y}\equiv\delta y-\frac{\partial f}{\partial({\rm d}\rho/{\rm d}p)}\delta\rho\,$, (75)
and we have initial condition $\delta y\big|_{p=p_{0}}=0$. Equation (74) now can be solved with a similar procedure we introduced in Sec. III. Considering $\delta\rho=\delta(p-p_{1})$, we can solve for $\delta\tilde{y}$ with
$\frac{{\rm d}\delta\tilde{y}}{{\rm d}p}=\frac{\partial f}{\partial r}\delta r+\frac{\partial f}{\partial m}\delta m+\frac{\partial f}{\partial y}\delta\tilde{y}\,$, (76)
and initial conditions
$p=p_{1}\,,\quad\delta\tilde{y}=-\left.\left(\frac{\partial f}{\partial\rho}-\frac{{\rm d}}{{\rm d}p}\frac{\partial f}{\partial({\rm d}\rho/{\rm d}p)}+\frac{\partial f}{\partial y}\frac{\partial f}{\partial({\rm d}\rho/{\rm d}p)}\right)\right|_{p=p_{1}}\,.$ (77)
We shall note that, one can directly integral Eq. (73) around $p_{1}$ to obatin the integration initial condition for $\delta y$ in $p=p_{1}^{-}$ that is equivalent to what we have done here. However, one should be careful that $\delta y$ contains a delta function part that also contributes to the jump condition, which is the last term in the above equation.
For calculating $\delta\Lambda$, we are interested in the value at the boundary, $\delta Y=\delta y\big|_{p=0^{-}}$, where we use $0^{-}$ to denote the point outside the star but infinitely close to the surface. One needs to be careful with the fact that, similarly to the calculation of the QS tidal Love number, the ${\rm d}\rho/{\rm d}p$ term in the equation leads to additional correction terms at the boundary. For example, the relevant term that is proportional to $\delta r$ reads
$\frac{{\rm d}\delta{y}}{{\rm d}p}=\cdots+\frac{12\pi r^{3}m}{(m+4\pi r^{3}p)^{2}}\frac{\delta r}{r}\frac{{\rm d}\rho}{{\rm d}p}+\cdots\,$, (78)
which leads to a correction term
$\delta{y}\big|_{p=0^{-}}-\delta{y}\big|_{p=0^{+}}=-\frac{12\pi R^{3}\rho_{s}}{M}\frac{\delta R}{R}+\cdots\,$, (79)
where $\rho_{s}$ is the density at the surface of the star and dots represent the corrections come from other terms. We have used $0^{+}$ to denote the point that is inside but infinitely close to the surface opposite to $0^{-}$.
With $\delta Y$ in hand, $\delta\Lambda$ then can be calculated via
$\delta\Lambda=\frac{\partial\Lambda}{\partial M}\delta M+\frac{\partial\Lambda}{\partial R}\delta R+\frac{\partial\Lambda}{\partial Y}\delta Y\,$, (80)
where $\Lambda$ given by Eq. (52) and Eq. (53) is regarded as a function of $M$, $R$, and $Y$.